OpenStudy (agentjamesbond007):

Prove that, for all x>1, 2√x > 3 - 1/x

OpenStudy (anonymous):

Hi

OpenStudy (agentjamesbond007):

Hello

OpenStudy (anonymous):

I think i might know how to solve this problem

OpenStudy (anonymous):

First we get all x's on one side

OpenStudy (agentjamesbond007):

I tried to isolate x to one side. I can't seem to find a way.

OpenStudy (anonymous):

now is the x under the 1 or is the x under the 3-1?

OpenStudy (agentjamesbond007):

the x is under the 1

OpenStudy (anonymous):

ok give me a moment to work it out on paper

OpenStudy (anonymous):

I think your statement about x>1 is false. For example,\[\frac{ 1 }{ 2 }\] is not greater than 1 but the equation is still true.

OpenStudy (agentjamesbond007):

So would that be a way to prove it?

OpenStudy (anonymous):

Prove it false yes.

OpenStudy (anonymous):

no you will need to get x by itself

OpenStudy (anonymous):

to actually prove it

OpenStudy (agentjamesbond007):

oh , ok.

OpenStudy (anonymous):

I don't know if it the sign was suppose to be written as ≥. If you plug in 1, you will get 2 > 2, which is false.

OpenStudy (agentjamesbond007):

I keep getting into a loop when I divide 2 to both sides

OpenStudy (anonymous):

@micahwood50 That does make sense. But it does not prove why \[\frac{ 1 }{ 2 }\] does work in the equation. Like i said i think the statement X>1 is false with the equation given.

OpenStudy (anonymous):

i think we need to get like terms

OpenStudy (anonymous):

Never mind, I didn't realize the statement "for all x>1"...

OpenStudy (anonymous):

@Agentjamesbond007 It seems as if you need to refine your statement "for all x>1" for us to solve it.

OpenStudy (agentjamesbond007):

Well, that is the prompt I'm needed to prove. I'm not sure how to rephrase it.

OpenStudy (anonymous):

\[1/x = x ^{-1}\] So using that you will have \[2√x > 3 -x ^{-1}\] You then will be able to square both sides and take the x's over.

OpenStudy (anonymous):

\[2\sqrt{x}>3-\frac{1}{x}\]We can assume \(x\) is positive, so we will multiply both sides by \(x\).\[2\sqrt{x^3}>3x-1\]We can then square both sides. Since we know \(x>1\), we know both sides are positive right now and greater than 1, so the equality will hold.\[4x^3>9x^2-6x+1\]\[4x^3-9x^2+6x-1>0\]Now we factor:\[(x-1)(x-1)(4x-1)>0\]As you can see, the only roots are \(1\) and \(\frac{1}{4}\). These are the only times it crosses, the x-axis and therefore the only times it can change sign. Since we know \(x>1\), we know \(x\) will never cross the x-axis. So if it is positive for say \(x=2\), then the inequality holds for all \(x>1\).\[((2)-1)((2)-1)(4(2)-1)>0\]\[(1)(1)(7)>0\]\[7>0\]